OpenStudy (anonymous):

-4(m + 6) = -8 wat is m? last question

OpenStudy (anonymous):

do you want the answer or explanation with answer?

OpenStudy (anonymous):

any is fine but explanation would be good

OpenStudy (anonymous):

you have to distribute the 4 to the m+6

OpenStudy (anonymous):

so..

OpenStudy (shamil98):

m = -4

OpenStudy (shamil98):

4(-4+6) = 8 4(2) = 8 hue

jigglypuff314 (jigglypuff314):

distributive property is a(b+c) = ab + ac so 4(m+6) = 4m + 24 so you now have 4m + 24 = -8 then subtract 24 from both sides -24 -24 4m = -32 then divide 4 by both sides /4 /4 so m = -32 divided by 4 = ?

OpenStudy (anonymous):



OpenStudy (anonymous):

-8 @jigglypuff314

OpenStudy (anonymous):

so 4 X m = 4m and 4 X 6=24, so you have 4m+24

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

wait its -4(m + 6) = -8

OpenStudy (anonymous):

4m+24=-8 then -24 on both sides 

OpenStudy (anonymous):

okay so -4m -24= -8

OpenStudy (shamil98):

you know.. there's a real easy way to do this